# SoK: “Interoperability vs Security” Arguments: A Technical Framework

### v1

@@ 3.1.1 Analysis: Security Engineering Concerns
1. How do interoperation and economic incentives clash (or not)? The incumbent companies currently have strong network effects that outweigh any incentives to interoperate. Apple can also bundle goods in a way that incentivizes exclusivity. The fact that iMessage is only available on iPhones and is not fully interoperable with messaging from Android (and other) phones incentivizes users to switch to iPhones. These incentives lead companies to simply avoid full interoperation, rather than to enact other punitive or exclusionary measures on competitors. Companies’ outsized power in this context is enough to put them in positions of market dominance.
@@
2. How do security arguments and economic incentives align (or not)? Interoperation is generally not economically advantageous for the companies being asked to interoperate. For WhatsApp, opening up a closed system is not only an engineering challenge; but it also means losing existing network effects. In the case of iMessage, extending functionality, including encryption, will be an engineering task and will remove an incentive for Android users to switch to iPhones. Here, however, there are also security arguments in the other direction: encrypting iMessage communications to/from Androids will better protect iMessage users. This is a interesting example of a company apparently eschewing security in favor of other economic incentives, yet at the same time framing some of their arguments for doing so as security arguments. This suggests that the utility and security of users is less important than exerting pressure on non-iPhone users to switch.
@@
3. What needs to change in order for the company to support interoperation? Although Meta already uses the Signal Protocol for WhatsApp, there is still a substantial amount of engineering to make sure that interoperation is secure. The Signal protocol is a centralized server design, the predominant paradigm in encrypted messaging, and Meta has also made some modifications to the Signal protocol in WhatsApp [51]. For Apple, some of the changes that are necessary for interoperation are simpler (such as the bubble color) whereas some are more complicated (such as encryption). In both cases, using community standards – such as Signal or Messaging Layer Security (MLS) – could achieve strong security for users as well as easier interoperation in the future.
@@
4. What tension exists between security and other goals both before and after? End-to-end encryption is a well-researched goal, and there are standards bodies and academic researchers as well as companies looking to improve the existing technology. There are open questions around how best to achieve interoperation in E2EE, especially between existing systems that need to be ‘retrofitted’ with interoperation capabilities [43]. At the same time, there are readily viable options using existing technology to set up interoperation that remains secure. Thus, security considerations are not an insurmountable obstacle to an interoperability mandate in this context; at the same time, research progress over the course of the next few years may reveal ways to do interoperable E2EE much better. In the case of iMessage, interoperating encryption for communications with Android users will ultimately better protect Apple’s own users. Thus, although initially difficult, security will be better served after the required changes.
@@ 3.2 Security Vetting Concerns
Our case studies in this section relate to Apple’s relationship to apps, so we begin with some background context. Apple has historically adhered to a ‘walled garden’ approach for iPhones. Users are only able to download and use apps vetted by Apple according to their security, privacy, and quality standards. The stringent screening of apps on the App Store has furthered the iPhone’s reputation as being more secure[41].
This approach limits what users can do on their own devices. Such limited usability appears to be untenable on PCs. [^13] Recall that we use “PC” broadly to include Macs; see footnote 7. Apple’s own approach to its Macs contrasts with its approach towards third-party software on its other devices. Apple has a notarization process for software intended for Macs, but this process is automated and much more limited than the process for apps on iOS [16, 8]. It is also cheaper—the cost to developers for Macs is also only the \$99 membership to the Apple Developer Program [9], which iOS developers pay in addition to other substantial fees [^14] See Case Study 3.2 and the Appendix for more on fees.—and easier—once notarized, this software can be distributed via project websites, not only on an Apple marketplace. These policies indicate Apple’s discretionary choices about the trade-off between user choice and security vetting on its products.
From the perspective of developers and antitrust officials, there is another dimension to this trade-off. Apple’s vetting of applications and its position as gatekeeper and security guarantor coincides with a position of market power. Apple’s control over its App Store allows it to determine what apps are offered and take a cut of transactions.
Thus, to cast security vetting concerns as only a balancing of user choice against security, is to miss the economic incentives at play in the background. The core reasons behind the push towards more open app distribution have been economic. Developers are not arguing for the freedom to distribute malware; rather they want to avoid the substantial rents that gatekeepers force upon them  [^15] Epic Games, Inc. v. Apple Inc., Case 4:20-cv-05640-YGR, Findings of Fact and Conclusions of Law Proposed by Epic Games, Inc. 2021. We now focus on four similar and overlapping examples of security vetting arguments. All have to do with app distribution and use.
@@
App sideloading refers to downloading an app onto a device via some route other than the official app store. This alternate route might be a website or alternative app store. Sideloading was not allowed on iPhones (or iPads) until the DMA, and is still not permitted outside the EU. Apple still requires that sideloaded apps be notarized, meaning the app is vetted and signed by Apple and only apps with such a signature can be downloaded onto a device. This notarization procedure is not as thorough as the vetting for the App Store, but still includes all the security requirements in the App Store Guidelines [8]. A major reason why Apple says they need to notarize all apps, especially those with alternative distribution, is to maintain the security standards of apps on its App Store [33].
The notarization process is not necessarily cheaper than getting approved for the App Store, despite there being fewer requirements for Apple to check. Developers that choose sideloading must pay a €0.50 ‘core technology fee’ (CTF) for each first annual install over one million, where annual installs include updates [13]. Apple allows free installs for a year by the same user, so within a year, the app can be updated multiple times for free, but this still means there is a charge to update each following year [21]. Updating apps to fix security issues is an important means of protecting users. Disincentivizing proactive updates runs counter to Apple’s stated goals of preserving security. Depending on a developer’s circumstances, this alternative pricing is not necessarily cheaper than the 30% cut Apple takes from the App Store and IAP; see Appendix for an example calculation of fees. Despite being designed to address concerns about Apple’s gatekeeing role and rent-seeking behavior, the notarization fee scheme has not reduced the burden of fees for all developers. Apple justifies these continued high fees by invoking user security.
@@
Apple must now allow alternative app stores, which will be able to distribute sideloaded apps. As discussed, the apps being sideloaded must be notarized by Apple and any in-app purchases would still pay the alternative payment commission (see Section 3.2).
The alternative app store itself must abide by additional criteria set by Apple, which Apple can check before allowing the alternative marketplace to operate. Alternative marketplaces must, for example, agree not to distribute apps that infringe on intellectual property rights or that are malicious as well as deal with requests for removal of apps by governments[12]. Apple also points out that collecting data about (insecure) apps through its App Store is an important aspect of keeping users secure [33]. There is no reason that interoperability regulations would prevent Apple from engaging in this type of data sharing with alternative app stores. Indeed, this type of data sharing would be beneficial even across existing marketplaces outside of Apple’s control, e.g., Google Play.
Thus far, these are vetting concerns, but there are also engineering considerations involved. Apple created an API to allow these alternative marketplaces to be able to execute downloads [12]. There is also minor amount of extra work that needs to be done for developers distributing their apps via their website[19]. Apple needs to set up these systems such that they interface with the OS correctly and securely. Though these are relatively lightweight engineering tasks, this provides a simple example illustrating a need for Apple to build secure architecture and vet third parties to interoperate.
@@
In-App Purchases (IAPs) are purchases of digital goods in apps on Apple devices, which are intermediated by Apple. We use IAP to refer specifically to Apple’s in-app purchase system. Apple generally takes a 30% cut from these purchases [20]. Payment for physical goods within apps is treated differently, but the purchase of in-app goods, like coins in a game, are required to go through Apple and incur the fee [8]. Securing payments is, of course important, and Apple’s justification for requiring that apps on its devices take payments through Apple’s system is to ensure the security of transactions[33]. [^16] It is unclear how this supports not enforcing intermediated payments for physical goods purchases.
Changes due to the DMA now allow developers to link out to websites or use alternative payment service providers for purchases of digital goods [17]. Processing payments online is simple and usually results in much smaller fees; Epic reports a figure less that 5% [^17] Epic Games, INC. v. Apple INC., Case 4:20-cv-05640-YGR, Findings of Fact and Conclusions of Law Proposed by Epic Games, Inc. 2021, §454. That said, developers who choose this option still have to pay a 10-17% commission on sales of digital goods and provide Apple with transaction reports [17].
From the perspective of the user, the new, open system is not dissimilar from the general state of online payments or the case for physical goods purchased on apps, where each merchant must independently assure the user of the safety of the transaction. Developers can use other widely trusted methods of payment or opt to use Apple’s IAPs. Thus, it falls to developers to vet their chosen payment service provider.
[…]
Prior to the DMA and the subsequent actions taken by the European Commission (henceforth ‘the Commission), all web browser apps on iOS had to use WebKit as their browser engine [5]. Browser engines are part of web browser apps (and in-app browsing) that render the content of a website into a page ready for user interaction [44]. WebKit is open source, but belongs to Apple[58]. As a result of Commission actions, Apple now allows alternative browser engines, only in the EU, both for browser apps and in-app browsing [6]. Prior to these changes, Apple’s iOS was the only operating system that did not support other browser engines; other browser engines can be used on the operating systems of their competitors [31]. In the developer documentation, the requirement that browsers use WebKit is not explicitly based on security, although there is notable community discussion around security motivations for this requirement [49].
@@ 3.2.1 Analysis: Security Vetting Concerns
“Security vs. interoperability” concerns in the vetting category exhibit a number of key patterns, as illustrated by the above case studies, and synthesized below.
@@
1. How do interoperation and economic incentives clash (or not)? The examples we have seen are similar in that Apple controls third-party developers’ access to their potential customer base on the iPhone. The iPhone is a platform and Apple can exclude developers on grounds that their apps do not meet Apple’s guidelines or because their services are not allowed in the ecosystem (e.g., alternative app stores). Apple’s monopolistic position also allows it to charge developers more than might be expected in a more competitive setting. This type of power is markedly different from the market power discussed in the security engineering case study and in horizontal interoperation more generally. Here, Apple has the power to completely exclude a third party, whereas, in horizontal interoperation, it cannot cleave third-party developers from users.
@@
2. How do security arguments and economic incentives align (or not)? Vetting apps to protect user security is a reasonable end, but it is also used as a means to extract rents. Keeping malware and other insecure apps away from users is also good for platforms, especially those that, like Apple, enjoy a reputation for security and good curation. That said, it is hard to evaluate whether the price tag is justified, given the opacity of the system and limited competition.
@@
3. What needs to change in order for the company to support interoperation? It is clear in all these cases that some amount of vetting is important to preserve the security status quo. Apple’s notarization system and CTF is Apple’s suggested way forward. The notarization guidelines are less onerous for Apple to check than their status quo, and vetting new app stores will likely be a rare occurrence—but the new mechanisms are not necessarily cheaper for developers. There have been changes to Apple’s pricing practices (and it is possible more changes will need to be made in the future). Price changes might be impactful for the company’s profits; they will not be difficult from an engineering standpoint.
@@
4. What tension exists between security and other goals both before and after? Apple has absolute control over third-parties’ access to iPhone users. At the same time, they rely on those third-parties to develop apps for the iPhone, increasing the iPhone’s utility and commercial success. The ecosystem of iOS apps is massive, and preventing every possible ‘bad’ outcome is too general and too far-reaching to be a viable reality. Instead, Apple has policies (e.g., [8]) around how they protect users and devices. Inherent in these policies are certain trade-offs around security, safety, privacy, and quality. Guidelines that are too stringent might unnecessarily narrow the universe of available apps or take too many resources. Guidelines that are too lenient might deeply undermine user trust and device integrity.
@@
The changes the Commission is enforcing will not necessarily impose new trade-offs when it comes to app vetting. Apple itself decided (perhaps for economic reasons) to apply only a subset of app guidelines to those apps applying for notarization. Alternative app payments and alternative app stores will require users to invest their trust in other third parties. Thus, arriving at an acceptable trade-off falls to the users; they must choose which companies’ security (and privacy) policies are trustworthy and/or worth the risk.
@@ 3.3 Hybrid Concerns
The hybrid cases are those in which the company must engineer a system that can preserve security when opened up to third parties, as well as vet those third parties.
@@
Near Field Communication (NFC) is a set of protocols that work at close range and allow two devices to communicate securely. NFC is used for myriad applications linking smartphones to other devices, including reading tags on objects, serving as keys, or communicating credit card or identification information [10]. For an app to use NFC for mobile payments, it must have access to the NFC antenna in the phone. It must also access key information that is stored on the Secure Element, which holds sensitive information, including the card information necessary for host card emulation (HCE) [7]. Apple had restricted third-parties from using NFC and the Secure Element for HCE, reserving this ability for Apple Wallet and Apple Pay. It did allow other usage of NFC via API access to the necessary hardware and software [14].
@@
EU antitrust authorities have identified this as monopolistic behavior in a case not based on the DMA [^18] CASE (EU) AT.40452 –Apple – Mobile Payments, Commission Decision, 2024 and Apple has opened up the HCE capability to third-party apps, subject to approval by Apple. The criteria for authorization require that the third-party developer commit to complying with laws and standards pertaining to privacy and security, including standard set by EMVCo, a standards body for card-based payments [14]. In addition to complying with existing standards, Apple requires that the third-party developer have certain policies around privacy and dealing with vulnerabilities in their product[14]. This is a vetting concern, and screening those developers looking to use HCE is an important aspect of protecting security in the new, open system.
@@
In order to ‘open up’ NFC capabilities for other wallet applications, third-party developers must have access to NFC, which is already available, as well as access to the Secure Element [11]. Access to the Secure Element is already separate from from the Secure Enclave, which is where the operating system stores other sensitive data [10]. This new access must be built correctly in order to preserve the security of both the third-party data and the rest of the system, and the precise architecture should be carefully designed to this end [^19] Android supports HCE both with and without a secure element, and allows third-party interoperation with the later method and developers cannot access the secure element [4]. Apple could potentially achieve interoperability without involving the secure element at all.. The Secure Element has internal boundaries that could impede a malicious developer from accessing other sensitive data and none of this data needs to come into contact with the operating system or app – the reply can go straight from the hardware to the payment processor [28]. Thus, APIs built for third-party developers could grant access to only specific functionalities without compromising others. There are also existing vulnerabilities [28], which would not necessarily be exacerbated by interoperability and might benefit from greater attention from the community. Thus, there is an engineering security concern that must be solved by Apple.
Since opening up this functionality to third-party developers requires significant engineering and vetting security concerns, we consider it a hybrid case under our framework.
@@
There are a number of other hardware and software features of iOS that the DMA has stipulated should be made available to third-party developers [^20] DMA.100203 – Article 6(7) – Apple – iOS – SP – Features for connected physical devices. As stated in the case, these measures relate to notifications on smartwatches, background execution, content casting and transfers, and paired device set up. Apple will be required to provide interoperability with the full list of functionalities to third parties. All of these functionalities are already available to Apple’s own devices (including Apple Watch, AirPods, and other iPhones), so achieving interoperability is a question of opening up previously reserved functionality to third-parties, generally via APIs and SDKs.
@@
Although the stipulations from the Commission are still evolving, Apple currently will evaluate interoperability requests on a case-by-case basis [15]. In their recent public-facing document, Apple mentions that Meta has made 15 requests for interoperation regarding various functionalities [15]. While Apple may be required to interoperate with competitors, the regulators do not require that any interoperation request be granted in a way that would undermine the security of the device [^21] REGULATION (EU) 2022/1925 on contestable and fair markets in the digital sector (Digital Markets Act), Article 6(7) . Thus, Apple could deny interoperation requests from malicious actors and create bespoke and secure solutions for interoperation to reduce security risk.
@@
One illustrative example is peer-to-peer (P2P) Wi-Fi connections, which can be used to transfer files, such as with Apple’s AirDrop feature, and continuity services, such as Handoff or Universal Clipboard[53]. These functionalities use a Bluetooth Low Energy (BLE) connection between two devices to create an ad hoc P2P Wi-Fi connection, which uses Apple’s iCloud credentials as part of its certificate procedure [10]. Several aspects of these services, including BLE, have been shown to leak user information or otherwise have vulnerabilities [25, 47, 37]. Apple Wireless Direct Link (AWDL) is an key protocol used in P2P Wi-Fi, which is based on IEEE 802.11, a set of standards, but is currently proprietary [54]. This protocol has been shown to have security vulnerabilities [54, 55, 53]. The fact that AWDL is proprietary is a barrier both to better study of its security and to interoperation [54].

### v2

@@ 4.1.1 Analysis: Security Engineering Concerns
1. How do interoperation and economic incentives clash (or not)? The incumbent companies currently appear to enjoy benefits from locking in users to their platforms (‘network effects’ [44]), which likely outweigh incentives to interoperate. [^19] This appears likely from the companies’ resistance to recent interoperability mandates; see also, e.g., [40] for an empirical perspective. Apple can also bundle goods in a way that incentivizes exclusivity. The fact that iMessage is only available on iPhones and is not fully interoperable with messaging from Android (and other) phones [18] may further incentivize users to switch to iPhones. These incentives may lead companies to simply avoid full interoperation, rather than to enact exclusionary measures on competitors.
@@
2. How do security arguments and economic incentives align (or not)? In many cases, interoperation may not be economically advantageous for the companies being asked to interoperate, as discussed above. For WhatsApp, opening up a closed system is not only an engineering challenge; it also means losing existing network effects. For iMessage, extensions to support full interoperation, including encryption, will be an engineering task and will remove an incentive for Android users to switch to iPhones. Here, however, there are also security arguments in the other direction: encrypting iMessage communications to/from Android better protects iMessage users. This is a interesting example of a company apparently eschewing security in favor of other economic incentives, yet at the same time framing some of their arguments for doing so as security arguments.
@@
3. What needs to change in order for the company to support interoperation? Although Meta’s messaging apps already build on the Signal Protocol for WhatsApp, there is still a substantial amount of engineering to make sure that interoperation is secure. The Signal protocol is a centralized server design, the predominant paradigm in encrypted messaging, and Meta has also made some modifications to the Signal protocol in WhatsApp [61]. For Apple, some of the changes that are necessary for interoperation are simpler (such as chat bubble color [18]) whereas some are more complicated (such as encryption). In both cases, using community standards—such as Signal or Messaging Layer Security (MLS)—could achieve strong security for users as well as easier interoperation in the future.
@@
4. What tension exists between security and other goals both before and after? End-to-end encryption is a well-researched goal, and there are standards bodies and academic researchers as well as companies looking to improve the existing technology. There are open questions around how best to achieve interoperation in E2EE, especially between existing systems that need to be ‘retrofitted’ with interoperation capabilities; at the same time, there are readily viable options using existing technology to set up interoperation that remains secure [49, 24]. Thus, security considerations are not an insurmountable obstacle to an interoperability mandate in this context; at the same time, research progress over the course of the next few years may reveal ways to do interoperable E2EE much better. In the case of iMessage, interoperating encryption for communications with Android users will ultimately better protect Apple’s own users. Thus, although initially difficult, security will be better served after the required changes.
@@ 4.2 Security Vetting Concerns
Next, we present three similar and overlapping relevant technologies, featuring distinct SvI concerns in the vetting category. All three case studies relate to app stores. We mostly focus on Apple; there have also been similar cases in the US against both Apple and Google. [^20] See Epic Games, Inc. V. Apple, Inc., No. 21-16506 (9th Cir. 2023) and Epic Games, Inc. v. Google LLC, No. 24-6274, (9th Cir. 2024)
We begin with some context. Apple has historically adhered to a ‘walled garden’ approach for iPhones. Users are only able to download and use apps vetted by Apple according to their security, privacy, and quality standards. The stringent screening of apps on the App Store has furthered the iPhone’s reputation for strong security[47].
This approach limits what users can do on their own mobile devices—far more strictly than on any existing PC systems. [^21] Recall that we use “PC” broadly to include Macs; see footnote 12. Apple’s own approach on its Macs contrasts with its approach towards third-party software on its other devices. Apple has a notarization process for software intended for Macs, but this process is automated and much more limited than the process for iOS apps [16, 8]. It is also cheaper—the cost to developers for Macs is membership in the Apple Developer Program (99 USD yearly) [9], which iOS developers must have in addition to other substantial fees. [^22] See the Appendix for more on fees. Once notarized, Mac software can be distributed via a website, not only via Apple marketplaces. These policies reflect Apple’s discretionary choices about tradeoffs between user choice and security vetting on different products.
From the perspective of developers and antitrust officials, there is another dimension to this tradeoff. Apple’s vetting of applications and its position as gatekeeper and security guarantor coincides with a position of market power. Apple’s control over its App Store allows it to determine what apps are offered and take a cut of transactions.
Thus, to cast security vetting concerns as only a balancing of user choice against security, is to miss the economic incentives at play in the background. The core reasons behind the push towards more open app distribution appear to have been economic. Developers are not arguing for the freedom to distribute malware; rather, they want to avoid the substantial rents that gatekeepers force upon them  [^23] Epic Games, Inc. v. Apple Inc., Case 4:20-cv-05640-YGR, Findings of Fact and Conclusions of Law Proposed by Epic Games, Inc. 2021.
@@
App sideloading refers to downloading an app onto a device via some route other than the official app store. This alternate route might be a website or alternative app store. Sideloading was not allowed on iPhones (or iPads) until the DMA, and is still not permitted outside the EU. Apple still requires that sideloaded apps be notarized, meaning the app is vetted and signed by Apple and only apps with such a signature can be downloaded onto a device. This notarization procedure is not as involved as the vetting for the App Store, but still includes all the security requirements in the App Store Guidelines [8]. A major reason why Apple says they need to notarize all apps, especially those with alternative distribution, is to maintain the security standards of apps on its App Store [32].
The notarization process is not necessarily cheaper than getting approved for the App Store, despite there being fewer requirements for Apple to check. Developers that choose sideloading must pay a 0.50 EUR ‘core technology fee’ (CTF) for each first annual install over one million, where annual installs include updates [13]. Apple allows free installs for a year by the same user, so within a year, the app can be updated multiple times for free; there is a charge to update each following year [21]. Updating apps to fix security issues is an important means of protecting users. As timely updates are important for security, disincentivizing proactive updates may undermine Apple’s stated goals of preserving security. Depending on a developer’s circumstances, this alternative pricing is not necessarily cheaper than the 30% cut Apple takes from the App Store and in-app purchases. [^24] See Appendix for an example calculation of fees. Apple justifies these fees by invoking the role it plays in developing tools for developers, which include security benefits from the notarization process [21, 22].
In addition to paying fees on purchases of apps, Apple requires all in-app purchases (IAP) of digital goods to be intermediated by Apple, and generally subjects these to a 30% fee [20] (this fee structure changed after the DMA in Europe, but remains in place elsewhere [^25] Even developers who choose a link out option now available in the EU still have to pay a 10-17% commission on sales of digital goods and provide Apple with transaction reports [17].). Outside the EU, Apple also forbids apps from linking to websites where purchases can be completed like any other online purchase (without Apple intermediating) [8]. Different rules govern the purchase of physical goods. The type of good (physical or digital) being purchased would appear to have no bearing on payment security. However, Apple’s justification for requiring usage of Apple’s IAP system is to ensure the security of transactions[32]. That said, intermediaries for processing payments securely online are common, and usually levy much smaller fees (e.g., Epic reports less than 5% [^26] Epic Games, INC. v. Apple INC., Case 4:20-cv-05640-YGR, Findings of Fact and Conclusions of Law Proposed by Epic Games, Inc. 2021, §454).
@@
Apple must now allow alternative app stores, which will be able to distribute sideloaded apps. As discussed above, the apps being sideloaded must be notarized by Apple.
The alternative app store itself must abide by additional criteria set by Apple, which Apple can check before allowing the alternative marketplace to operate. Alternative marketplaces must, for example, agree not to distribute apps that infringe intellectual property rights or that are malicious as well as deal with requests for removal of apps by governments[12]. Apple also points out that collecting data about (the security of) apps through its App Store is an important aspect of keeping users secure [32]. There is no reason that interoperability regulations would prevent Apple from engaging in or requiring this type of data sharing with alternative app stores [32]. Indeed, this type of data sharing would be beneficial even across existing marketplaces outside of Apple’s control, e.g., Google Play.
The vetting concerns have been the prominent ones in these cases, but there are also engineering considerations involved. Apple created an API to allow these alternative marketplaces to be able to execute downloads [12]. There is also minor amount of extra work that needs to be done for developers distributing their apps via their website[19].
@@
Prior to the DMA and the subsequent actions taken by the Commission, all web browser apps on iOS had to use WebKit as their browser engine [5]. Browser engines are part of web browser apps (and in-app browsing) that render the content of a website into a page ready for user interaction [51]. WebKit is open source, but belongs to Apple[71]. As a result of Commission actions, Apple now allows alternative browser engines, only in the EU, both for browser apps and in-app browsing [6]. Prior to these changes, Apple’s iOS was the only operating system that did not support other browser engines; other browser engines can be used on the operating systems of their competitors [68]. In the developer documentation, the requirement that browsers use WebKit is not explicitly based on security, although there is notable community discussion around security motivations for this requirement [59].
@@ 4.2.1 Analysis: Security Vetting Concerns
1. How do interoperation and economic incentives clash (or not)? The three case studies share the important context of Apple’s control over third-party developers’ access to their potential customer base on the iPhone. iPhone and iOS comprise a relatively closed device and platform ecosystem, and Apple can exclude developers on grounds that their apps do not meet Apple’s guidelines or because their services are not allowed in the ecosystem (e.g., alternative app stores). Apple’s monopolistic position definitionally also allows it to charge developers more than would be expected in more competitive settings. This type of ‘gatekeeper’ power differs markedly from the market power discussed in the security engineering case study above, and in horizontal interoperation more generally. Here, Apple has the power to completely exclude a third party; in horizontal interoperation, it cannot cleave third-party developers from users.
@@
2. How do security arguments and economic incentives align (or not)? Vetting apps to protect user security is a reasonable end, but it is also used as a means to extract rents. Keeping malware and other insecure apps away from users is also good for platforms, especially those that, like Apple, enjoy a reputation for security and good curation [47]. It is hard to (externally) evaluate whether the price tag is justified, given system opacity and limited competition.
@@
3. What needs to change in order for the company to support interoperation? In all these cases, some amount of vetting is important to preserve the security status quo. Apple’s notarization system and CTF is Apple’s suggested way forward. As discussed above, the notarization guidelines are less onerous for Apple to check than their status quo, and vetting new app stores will likely be a much rarer occurrence than vetting new apps—but Apple’s new mechanisms are not necessarily cheaper for developers. There have been changes to Apple’s pricing practices (and it is possible more changes will need to be made in the future). Price changes might be impactful for the company’s profits, but changing pricing is typically not an engineering challenge.
@@
4. What tension exists between security and other goals both before and after? Apple has absolute control over third-party developer access to iPhone users. At the same time, they rely on those third parties to develop apps for the iPhone, increasing the iPhone’s utility and commercial success. The ecosystem of iOS apps is massive, and preventing every possible ‘bad’ outcome is too general and too far-reaching to be a viable reality. Instead, Apple has policies (e.g., [8]) around how they protect users and devices ‘well enough.’ Inherent in these policies are tradeoffs around security, safety, privacy, and quality. Too-stringent guidelines might unnecessarily limit available apps or take too many resources. Too-lenient guidelines might undermine user trust and device integrity.
@@
The changes the Commission is enforcing will not necessarily impose new tradeoffs when it comes to app vetting. Apple itself decided to apply only a subset of app guidelines to those apps applying for notarization (whether for economic reasons or otherwise). Alternative app payments and alternative app stores will require users to invest their trust in other third parties. Then, choosing an acceptable tradeoff falls in larger part to the users; they must choose which companies’ security (and privacy) policies are trustworthy and/or worth the risk.
@@ 4.3 Hybrid Concerns
Recall that hybrid cases are those in which the company must engineer a system that can preserve security when opened up to third parties, as well as vet those third parties.
@@
We identified Near Field Communication (NFC) as a relevant technology from both the Apple Mobile Payments antitrust case  [^27] CASE (EU) AT.40452 Apple - Mobile Payments and the Qualcomm/NXP Semiconductors merger [^28] CASE (EU) M.8306-QUALCOMM / NXP SEMICONDUCTORS. The Qualcomm/NXP Semiconductors merger mandated interoperation between NFC chips and other chips whereas the Apple case mandates interoperation between those types of chipsets and software. (There is also a similar NFC-related antitrust proceeding in the US against Apple. [^29] United States v. Apple Inc. 2:24-cv-04055 (JXN-LDW) (D.N.J.) ) In our search for SvI concerns, we did not find any articulated concerns regarding the merger so we focus on the Apple case. The ‘security vs. interoperability’ considerations discussed by Apple include that new APIs had to be built to allow third-party developers secure access to the technology and that those developers will be vetted before such functionality is available [11], which falls into our hybrid definition.
@@
NFC is a set of protocols that work at close range and allow two devices to communicate securely. NFC is used for myriad applications linking smartphones to other devices, including reading tags on objects, serving as keys, or communicating credit card or identification information [10]. For an app to use NFC for mobile payments, it must have access to the NFC antenna in the phone. It must also access key information that is stored on the Secure Element, which holds sensitive information, including the card information necessary for host card emulation (HCE) [7]. Apple had restricted third parties from using NFC and the Secure Element for HCE, reserving this ability for Apple Wallet and Apple Pay. It did allow other usage of NFC via API access to the necessary hardware and software [14].
@@
EU antitrust authorities identified this as monopolistic behavior and Apple has opened up the HCE capability to third-party apps, subject to approval by Apple. The approval criteria require third-party developers to commit to complying with laws and standards related to privacy and security, including standards set by EMVCo, a standards body for card-based payments [14]. In addition, Apple requires third-party developers to have certain policies around privacy and dealing with vulnerabilities [14]. This is a vetting concern, and screening developers seeking to use HCE is an important aspect of protecting security in the newly open system.
@@
In order to ‘open up’ NFC capabilities for third-party wallet applications, third-party developers must have access to NFC, which is already available, as well as access to the Secure Element [11]. [^30] Android supports HCE both with and without a secure element, and allows third-party interoperation with the latter method; third-party Android developers cannot access the secure element [4]. If Apple adopted a similar approach, it could potentially achieve interoperability without involving the secure element at all; it has not opted for this approach. Access to the Secure Element is already separate from from the Secure Enclave, which is where the operating system stores other sensitive data [10]. This new access must be built correctly in order to preserve the security of both the third-party data and the rest of the system. The Secure Element has internal security boundaries that could impede a malicious developer from accessing other sensitive data and none of this data needs to come into contact with the operating system or app—the reply can go straight from the hardware to the payment processor [28]. Thus, APIs built for third-party developers could grant access to only specific functionalities without compromising others. There are also existing vulnerabilities in these products [28], which would not necessarily be exacerbated by interoperability and might benefit from greater attention from the community. Thus, there are a range of security engineering security concerns to be solved by Apple.
@@
We identified OS and physical device interoperation as a type of relevant technology from three cases: two similar cases involving Apple’s iOS [^31] CASE (EU) DMA.100204 SP - Apple - Article 6(7) – process and CASE (EU) DMA.100203 Apple – Operating systems – iOS – Article 6(7) – SP – Features for Connected Physical Devices and the Google/Fitbit merger [^32] CASE (EU) M.9660 GOOGLE/FITBIT. The Apple cases concern the interoperability between many aspects of iOS and third-party hardware and software whereas the merger focuses specifically on the interoperation between third-party hardware devices and the Android OS. (The US case concerning NFC, mentioned above, [^33] United States v. Apple Inc. 2:24-cv-04055 (JXN-LDW) (D.N.J.) also discusses physical device interoperation in a similar light to the EU cases about Apple.) The SvI arguments regarding the Apple cases are exemplified in a whitepaper by Apple [15].
Public discussions about security in the Google/Fitbit merger case were notably very different; they did not discuss the security of the mandated interoperability, but rather discussed distinct voluntary interoperability that would result from the merger, reassuring users that it would not endanger user security or privacy [58]. This type of argument is outside our current scope, so we focus on the Apple case for the rest of the case study. However, investigating this type of ‘inverted’ case—and examining how perspectives on the the compatibility of security and interoperability feature differently across economic contexts with different incentives—could make for interesting future work.
@@
The Apple cases cover many different hardware and software features of iOS, including notifications on smartwatches, background execution, content casting and transfers, and paired device set up. Apple will be required to provide interoperability with the full list of functionalities to third parties. All of these functionalities are already available to Apple’s own devices (including Apple Watch, AirPods, and other iPhones), so achieving interoperability is a question of opening up previously reserved functionality to third parties, generally via APIs and SDKs.
@@
Although the stipulations from the Commission are still evolving, Apple is currently evaluating interoperability requests on a case-by-case basis [15]. In their recent public-facing document, Apple mentions that Meta has made 15 requests for interoperation regarding various functionalities [15]. While Apple may be required to interoperate with competitors, the regulators do not require that any interoperation request be granted in a way that would undermine the security of the device. [^34] REGULATION (EU) 2022/1925 on contestable and fair markets in the digital sector (Digital Markets Act), Article 6(7) Thus, Apple could deny interoperation requests from known malicious actors and develop/require bespoke solutions for interoperation to reduce security risk.
@@
One example is peer-to-peer (P2P) Wi-Fi connections, which can be used to transfer files, such as with Apple’s AirDrop, and continuity services, such as Handoff or Universal Clipboard[65]. These functionalities use a Bluetooth Low Energy (BLE) connection between two devices to create an ad hoc P2P Wi-Fi connection, which uses Apple’s iCloud credentials as part of its certificate procedure [10]. Several aspects of these services, including BLE, have been shown to leak user information or otherwise have vulnerabilities [26, 55, 42]. Apple Wireless Direct Link (AWDL) is an key protocol used in P2P Wi-Fi, which is based on IEEE 802.11, a set of standards, but is currently proprietary [66]. This protocol has been shown to have security vulnerabilities [66, 67, 65]. AWDL being proprietary is a barrier both to better study of its security and to interoperation [66].